How to solve this
(x+4 upon 3) (x+1 upon 3)

Grade:8

1 Answers

snehanshu
8 years ago
(x+4)/3*(x+1)/3
=1/3.{(x+4)(x+1)}
using identity (x+a)(x+b)=x^2+x(a+b)+ab, we get
=1/3.{x^2+x(4+1)+4*1}
=1/3.{x^2+5x+4}
={x(x+5)+4}/3}         ,if you are unable to understand any part of this answer please do specify I would try my best to make you understand